Gravitational due to an infinitely long thin rod of linear density P from a point of mass M and perpendicular distance from rod to point is L

Grade:11

1 Answers

Sai Ram Charan
6 years ago
Data provided is inconsistent; anyway, we have the following data-
Length of rod=1/0 , magnitude of point mass= ‘m’​, distance of seperation= l
Therefore, from Newton’s Law of universal Gravitation, we have F=
 
-G{m_{1}m_{2} \over {\vert \mathbf {r} _{12}\vert }^{2}}\,\mathbf {\hat {r}} _{12}}(in vector calculuss form)
Clearly, the force is
 
 \frac{G P \frac{1}{0} m}{l^{2}}